OpenStudy (anonymous):

What are the intercepts of the line X-3y=9 A X-intercept:-3; Y-intercept:-9 B X-intercept:3; Y-intercept:9 C X-intercept:-9; Y-intercept:-3 D X-intercept:9; Y-intercept:-3

OpenStudy (anonymous):

just plug 0 in for x then solve to get y intercept and plug in 0 for y to get x intercept

OpenStudy (anonymous):

x - 3y = 9 (0) - 3y = 9 solve for y u get y-intercept

OpenStudy (anonymous):

yes u are correct

OpenStudy (anonymous):

what do i do next

OpenStudy (anonymous):

no i thought u meant it's the answer 9 and -3........ solveing for y u should have y = 9/-3

OpenStudy (anonymous):

thats -3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yup because ur saying y = -3 when x = 0 that is where the graph will cross the y-axis (0,-3)

OpenStudy (anonymous):

u want to find where the graph crosses the x axis by setting y = 0 and solve for x

OpenStudy (anonymous):

x - 3y = 9 x -3(0) = 9 solve for x

OpenStudy (anonymous):

your welcome i hope ur able to understand it better

OpenStudy (anonymous):

well what answer did u get for x?

OpenStudy (anonymous):

-3(0) = 0 so x = 9

OpenStudy (anonymous):

oh so is D

OpenStudy (anonymous):

yes good job

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):

it might help if u do this problem again later or tomorrow .... practicing will help just a suggestion. yeah your right i'm am pursuing to become a math teacher.

OpenStudy (anonymous):

your welcome glad i could help

OpenStudy (anonymous):

ty very much

OpenStudy (anonymous):

i'll try to help, i don't know everything lol

OpenStudy (anonymous):

do i just plug in the numbers and see what works

OpenStudy (anonymous):

think y intercept solve for y when x = 0......... yes.....think x intercept and solve for x when y = 0.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

you could do it and plug those numbers from the multiple choice but u'd have to do them separate and if they = each other then that would be one of the intercepts then do the other.......u could think of it x= 0 then y intercept could be plugged into the equation...either way u should get same answer. there are usually a lot of ways to solve problems.

OpenStudy (anonymous):

whatever way u feel comfortable with you should use. doesn't hurt to know more than one way though

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

u get the answer?

OpenStudy (anonymous):

u understand how to set it up ? y intercept is when x = 0 cuz the coordinate will be (0,y) on they axis.... same for x just reverse

OpenStudy (anonymous):

so answer would be (3,0) and (0,-2)......... the set up u could say 6x -9y =18 for y intercept use x = 0 and 6x-9y =18 for x intercept use y = 0 6(0) - 9y =18 6x-9(0)=18

OpenStudy (anonymous):

the other way to plug in numbers would be 6(-3) -9(2) = 18 set both to 2 equations like 6(-3) = 18 and -9(2) = 18 u find out neither equal 18 so u'd try next answer. this strategy might take longer though

OpenStudy (anonymous):

glad u got it

OpenStudy (anonymous):

nice keep up the good work. education is such a vital asset so keep striving. your welcome

OpenStudy (anonymous):

cool yw take care

OpenStudy (anonymous):

ur asking me y i helped u?

OpenStudy (anonymous):

thanks again for all your help

OpenStudy (anonymous):

no problem take care

OpenStudy (anonymous):

ok bye